The pressure 55.0 meters under water is 641 kPa.

Conversion factors 1 kPa=9.869×10−3 atm 1 atm=760 mmHg

What is this pressure in atm? What is this pressure in mmHg?
Answer #1

o 1 kPa =9.869x10 3 atm 0 = 641 kPa x 9.8698163 = 632 6-029 x103 = 6-326 atm 1kpa = 7.500 = 641x7.500 =480765 mmig . =4.807 1